The diameters of apples from a certain farm follow the normal distribution with mean 4 inches and standard deviation 0.4 inch. Apples can be size-sorted by being made to roll over mesh screens. First the apples are rolled over a screen with mesh size 3.5 inches. This separates out all the apples with diameters less than 3.5 inches. Second, the remaining apples are rolled over a screen with mesh size 4.3 inches. Find the proportion of apples with diameters less than 3.5 inches.

  • Round your answer to four decimal places.
Answer #1

Solution:

Given: The diameters of apples from a certain farm follow the normal distribution with mean 4 inches and standard deviation 0.4 inch.

Thus X ~ Normal (u = 4,0 = 0.4)

We have to find: the proportion of apples with diameters less than 3.5 inches.

P( X < 3.5 )= ...............?

Find z score for x = 3.5

1 -

3.5-4 -0.5 0.4 -= -1.25

Thus we get:

P( X < 3.5 )= P( Z < -1.25 )

Look in z table for z = -1.2 and 0.05 and find corresponding area.

.07 .08 .09 | 1: -3.4 -3.3 -3.2 -3.1 -3.0 -2.9 -2.8 -2.7 -2.5 -2.5 -2.4 -2.3 -2. 2 -2.1 -2.0 -19 -1.8 -1.7 -1.6 -1.5 -1.4 -13

P( Z < -1.25 )= 0.1056

thus

P( X < 3.5 )= P( Z < -1.25 )

P( X < 3.5 )= 0.1056

Thus the proportion of apples with diameters less than 3.5 inches is 0.1056